OpenStudy (anonymous):

b+6 divided by 5 = 10 HElP

OpenStudy (anonymous):

What is divided by 5 = 10? 50! x + 6 = answer :D

OpenStudy (anonymous):

*What divided by 5 = 10 Sorry I just woke up haha

OpenStudy (anonymous):

ha umm let me check hold on :)

OpenStudy (anonymous):

o snap i think its b+6x5=10 sorry

OpenStudy (anonymous):

Is x=times or does x= a variable?

OpenStudy (anonymous):

times

OpenStudy (anonymous):

ah okay so b + 30 = 10 what is 10 - 30?

OpenStudy (anonymous):

ha 20 ??

OpenStudy (anonymous):

lol no xD -20 silly

OpenStudy (anonymous):

then -20 + 30 = 10

OpenStudy (anonymous):

o yeah i new that :)

OpenStudy (anonymous):

lol no worries. Good luck!

OpenStudy (anonymous):

you just earn yourself a new fan

OpenStudy (anonymous):

holla

OpenStudy (anonymous):

>_< ok lol \